OpenStudy (anonymous):

Determine whether the equation defines y as a function of x. Show work: xy + 6y = 1

OpenStudy (princeharryyy):

y*(x+6) = 1 y= 1/(x+6)

OpenStudy (princeharryyy):

yes it defines..))

OpenStudy (anonymous):

y*(x+6) = 1 y= 1/(x+6)

OpenStudy (princeharryyy):

@kellyconnelly